3 ducks and 2 ducklings weigh 32 kg. 4 ducks and 3 ducklings weigh 44 kg. all ducks weigh the same as each other and all ducklings weigh the same as each other. What is the weight of 2 ducks and 1 duckling?
Can you please help me solve this. Thank you.

Answer by TimothyLamb(4379)

x = kg weight of a duck
y = kg weight of a duckling
---
3x + 2y = 32
4x + 3y = 44
---
put the system of linear equations into standard form
---
3x + 2y = 32
4x + 3y = 44
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = kg weight of a duck = 8
y = kg weight of a duckling = 4
---
answer:
2x + 1y = ???
2*8 + 1*4 = 20 kg
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---